OpenStudy (anonymous):

20. The mass of an elevator cabin is M. While pending to move upward the tension in the supporting wire will be twice to that while pending to move downward. The corresponding accelerations will be equal in magnitude but of opposite direction. Find the magnitude of acceleration of the elevator cabin while pending to motion. The acceleration of gravity is g = 9.81m/ s /s.

OpenStudy (mani_jha):

When going downward, \[T1=mg+ma\] While going upward, \[T2+ma=mg\] \[T2=ma-mg\] T2=2T1 Put that, and you will get a.